In a scale drawing of a house, the living room is 1 in long. The actual length is 24 ft. On the drawing, a hallway is 1/4 in. wide. Find the actual width.
Answer by unlockmath(1688)

Hello,
We can set this up as a proportion as follows:
1/24=(1/4)/x Now, multiply each side by x to give us:
(1/24)x=1/4 Then multiply each side by 24 will result in:
x=6 There we go. The hallway is 6 feet wide.
Make sense?
RJ
Check out a new book I wrote:
www.math-unlock.com